OpenStudy (anonymous):

Find the equation of the tangent line to the curve y=6tan(x) at point (pi/4, 6)

OpenStudy (anonymous):

thanks @suvesh253 ! can u help me one more time with this one?!

OpenStudy (calculusfunctions):

I'm assuming you know calculus?

OpenStudy (anonymous):

yes!

OpenStudy (calculusfunctions):

OK, so then you know that you have to first find the derivative (slope function), correct?

OpenStudy (calculusfunctions):

In order to find the equation of the tangent line, we must first find the slope of the tangent line, correct?

OpenStudy (calculusfunctions):

That is incorrect because\[\frac{ d(\tan x) }{ dx }=\sec ^{2}x\]

OpenStudy (calculusfunctions):

Please try again.

OpenStudy (calculusfunctions):

If\[y =6\tan x\]then \[\frac{ dy }{ dx }=?\]

OpenStudy (anonymous):

no clue

OpenStudy (calculusfunctions):

Derivatives of primary trigonometric functions: The derivative of sine is cosine. The derivative of cosine is negative sine. The derivative of tangent is secant squared. Knowing this, to take the derivative of any primary trigonometric function: i). take the derivative of the angle and multiply by the amplitude (the number in front of the trig function). ii). multiply this by the derivative of the trig function and always keep the angle the same. Now try again.

OpenStudy (calculusfunctions):

For example:\[y =3\sin (2x + 3)\] \[\frac{ dy }{ dx }=3[\frac{ d(2x +3) }{ dx }](\cos (2x +3))\] \[\frac{ dy }{ dx }=(3)(2)\cos (2x +3)\] \[\frac{ dy }{ dx }=6\cos (2x +3)\]Do you understand?

OpenStudy (calculusfunctions):

@tenistaego, where did you go?